How do I set up and solve: A ladder rests against ahouse at 15ft, the length of the ladder is 5 feet more than the distance from the house. What is the length of the ladder?
Answer by rfer(16322)

15^2+5^2=c^2
225+25=250
sqrt 250=15.8 ft